) If the moment generating function (m.g.f.) of a random variable X is 

( ) exp 3( 32 ). 2 M t t t X = + Find mean and standard derivation of X and also compute 

P(x < )


1
Expert's answer
2021-04-25T15:43:52-0400

Given mgf is-

"f(t)=e^{3t}"


First moment- Mean is given by


"E(X)=\int_0^1te^{3t}dt=t\dfrac{e^{3t}}{3}-\int_0^1\dfrac{e^{3t}}{3}dt"


"=\dfrac{2}{9}e^3-\dfrac{1}{9}"

Second moment -

"E(X^2)=\int_0^1t^2e^{3t}dt=t^2\dfrac{e^{3t}}{3}-\int_0^1\dfrac{2}{3}te^{3t}dt"

"=\dfrac{5}{27}e^3+\dfrac{2}{27}"


Standard deviation "=\sqrt{E(X^2)-[E(X)]^2}"


"=\sqrt{\dfrac{5}{27}e^3+\dfrac{2}{27}-\dfrac{4}{81}e^6-\dfrac{1}{81}+\dfrac{4}{81}e^3}"


"=\sqrt{\dfrac{19}{81}e^3-\dfrac{4}{81}e^6+\dfrac{5}{81}}"



P(X<a) can be computed by integrating f(t) from 0 to a.

"P(X<a)=\int_0^ae^{3t}dt=e^{3t}|_0^a=e^{3a}-1"